### v2

@@ Abstract
Electronic Health Records (EHRs) frequently contain binary event sequences, such as diagnosis, procedure, and medication histories. These sequences, however, are characterized by a large percentage of zero values, presenting significant challenges for model learning due to the scarcity of gradient updates during processing. We propose a novel approach that reframes EHR data from a statistical perspective, treating it as a sample drawn from patient cohorts and transforming binary events into continuous Bernoulli probabilities. This transformation not only models deterministic binary events as distributions but also captures event-event relationships through conditional probabilities. Building upon this probabilistic foundation, we implement a graph neural network that learns to capture complex event-event correlations while emphasizing patient-specific features. Our framework is evaluated on the task of medication recommendation, demonstrating state-of-the-art performance on large-scale databases. Extensive experimental results show that our proposed BernGraph achieves 0.589, 0.846, and 0.844 of Jaccard, F1 score, and PRAUC, respectively, on the MIMIC-III dataset, with improvements of 4.6%, 6.7%, and 6.3% compared to baselines, including those using secondary information. This work presents a promising direction for handling sparse binary sequences in EHRs, potentially improving various healthcare prediction tasks. The source code is available at https://github.com/chenzRG/BEHRMecom
@@ 1 Introduction
Figure 1: Illustration of medication recommendation systems. (a) Such systems can provide medication recommendations based on historical medical events. (b) The vast majority of event data are zero values. For example, the proportion of non-zero values exceeding 10% is lower than 0.9% in MIMIC-III. (c) Those zero values pose a challenge to model learning since they do not propagate learning signals.
@@
Medication errors are one of the most serious medical errors that could threaten patients’ lives. Studies indicate that over 42% of these errors stem from physicians or doctors with insufficient experience or knowledge concerning specific drugs and diseases [28]. This challenge becomes even more pronounced with more new drugs and treatment guidelines. On the other hand, many patients are concurrently diagnosed with multiple diseases, complicating the decision-making process for doctors when determining more appropriate remedies for patients [33]. To this end, the medical community will benefit from an automated medication recommendation system that can assist doctors in decision-making. In the past decades, a vast amount of clinical data representing patient health status (e.g., medical reports, radiology images, and allergies) has been collected. This has remarkably increased digital information, known as electronic health records (EHR), available for patient-oriented decision-making. In this context, mining EHR data for medication recommendation has attracted growing research interests [33], as illustrated in Fig. 1(a).
@@
Successful methods typically rely on modeling outcomes of various medical events and diagnoses by observing longitudinal patient histories, i.e., tracking state variations across hospital visits [3, 38, 36, 33]. They often employ sequential neural networks, such as RNNs or Transformers. They construct a feature space centered around visits, each characterized by the current state of various medical events. Their objective is to discern long-range temporal dependencies that reflect the historical progression of diseases. These methods, however, suffer from a number of drawbacks. From a data viewpoint, different EHR databases adopt distinct visit recording methodologies [24]. For example, the AMR-URI database [14] employs a regular visit record, while visits in the MIMIC database [10] are patient-specific. Even within the same database, the time intervals between visits can differ greatly between patients. From a model perspective, a common obstacle is the "cold-start" issue, where medication recommendation systems struggle due to insufficient patient-specific historical data [37]. In practice, this issue becomes particularly pronounced when the system encounters newly onboarded patients with scarce or even absent historical data (i.e., visits) about their medical backgrounds. While many researchers treat irregular time intervals as input variables [18, 15, 34] or apply the carry-forward imputation method [13], these approaches may introduce biases to temporal dependencies.
@@
Some studies focus solely on a patient’s current diagnoses and medical event records, modeling event-event correlations for recommendation purposes without considering historical visits [38, 25]. However, a performance gap remains when compared to visit-based methods. A typical medical event has a binary value indicating its outcome: 1 for occurrence and 0 for non-occurrence, which leads to a sparsity issue in EHR data. The sparsity, characterized by a high proportion of zero entries, can significantly impact the performance of recommendations (seen in Fig. 1(b)). Zero values provide no tangible information for deep learning models, preventing the activation of weight propagation. This often leads to a vanishing gradient problem and yields collapsed representations for medication recommendations (as shown in Fig. 1(c)). Previous studies employ secondary information, e.g., Drug-Drug Interaction Knowledge Graph (DDI-KG), to complement information to medical event features [33]. Yet, sparsity remains a persistent issue.
@@
Therefore, this paper presents BernGraph, a binary EHR data-oriented drug recommendation system. Our method is built upon a graph neural network (GNN) and solely utilizes the $\mathbb{EHR}$ binary medical events for recommendation. To tackle the binary data/sparsity, we take a statistical perspective that allows us to transform the 0, 1 values into continuous probabilistic representations, and estimate them from patient cohorts. Our statistical strategy has demonstrated substantial efficacy in modeling graphical structures and correlations within binary EHR data. Architecturally, diverse events are incorporated as node attributes within the GNN. When two events co-occur, we model their relationship as conditional Bernoulli probabilities to initialize the edge between them. This allows GNN to learn the correlations between events across different patients for medication recommendations. Our results demonstrate that our method outperforms several benchmarks, including the latest state-of-the-art (SOTA) visit-based methods. Specifically, our method shows an improvement on the MIMIC III dataset of approximately 4.8%, 6.7%, 6.3%, and 4% in terms of Jaccard, F1-score, PRAUC, and AUROC, respectively. BernGraph is a simple and effective model that complements existing research in which the EHR data is often exploited in conjunction with secondary data but without transformations.
@@ 2 Related Works
- EHR-based Medication Recommendation. Historical Visit-based Methods: Many studies use deep learning to extract features from state variations across hospital visits in EHR data [5, 31, 30, 2, 25]. For example, DoctorAI [5] employs RNN to learn the dependencies of patients’ historical information and the doctor’s medication recommendations record for recommendation. MICRON [35] focuses on historical medication changes and medication combinations of the last visit. Some works introduce additional information to assist drug prediction, e.g., Drug-Drug Interaction (DDI) [36, 33].
Current Medical Event-based Methods: On the other hand, some works aim to analyze the patient’s current health status: recommendations are made based on the EHR records from each patient visit [38, 9, 23]. LEAP [38] encodes patient information in a one-hot encoding manner and then learns the relationship between patient features and drugs. SMR [9] incorporates additional information. It combines EHR data with an external medical knowledge graph to leverage external medical expertise in aiding the model decision-making process. MT-GIN [23] models event-event correlations in EHR data using a graph-based approach, with a focus on events that co-occur. This approach overlooks the information that 0 values in EHR binary data can potentially provide.
- Graph Modeling for Medication Recommendation. Recent research considers that event-event correlations can be well-captured by graph structures [5, 4, 25, 32, 22, 9, 27, 33, 23]. For instance, Gamenet [22] and Cognet [33] employ GNNs to discern correlations between different data types. They integrate DDI data into GNNs to understand the relationship graph between drugs, facilitating representation learning for medical events. GATE [25] proposes a GNN medication recommendation using events collected during a single visit. Further, COGNet [33] simultaneously models the EHR graph and the DDI graph, subsequently integrating them into a generative model. However, the above methods are designed on top of the binary value of medical events.
- Binary Data Representation. Mining binary data has been a longstanding challenge [7, 29]. One of the reasons is, that 0s and 1s can have entirely different meanings in various contexts or projects, making it difficult to interpret them uniformly [29]. Moreover, 0 values often provide nothing to the learning process, especially in deep learning models. In the EHR context, medGAN and POPCORN explicitly transform binary events in the EHR dataset into continuous probability values [6, 2]. However, these efforts primarily focus on modeling historical information or learning an overall distribution of the medical event outcomes. An effective medication recommendation method that directly represents binary events in EHR data remains challenging.
@@ 3.1 Recommendation System Based on EHR
EHR data consists of various medical events, such as multiple symptoms, recorded chronically from patients. Each data sample contains binary outcomes of these events. Let $\boldsymbol{x}\in\{0,1\}^{M}$ denote a patient EHR sample, where $M$ represents the total number of medical events, $1$ stands for the occurrence of the corresponding event, and 00 for the absence. The collection of EHRs is a matrix $\mathbf{X}\in\{0,1\}^{N\times M}$ composed by stacking $N$ patient vectors $\{\boldsymbol{x}\}_{1:N}$.
It is especially difficult to build effective recommendation systems using solely the EHR data, mainly due to the following reasons: (1) binary data poses a challenge to effective machine learning: while both 0 and 1 play an equally important role in identifying a proper recommendation, in practice, 0 entries generally do not propagate learning signals; (2) medical events differ across platforms, and as a result, two binary matrices may share the same shape but have completely different underlying meanings. Moreover, extracting such meanings and properly representing them, be it a series of same events recorded chronically, or different events that are highly correlated, is nontrivial.
Our goal is to effectively mine the complete information hidden in binary EHR data, based on which an effective, general-purpose recommendation system can be built. Formally, BernGraph can be described by $F(\cdot):\{0,1\}^{M}\rightarrow\{0,1\}^{C}$, which accepts an EHR sample as input, and outputs $C$ binary indicators indicating whether or not recommend drug $i\in\{1,\dots,C\}$.
@@ 3.2 EHR as a Sample from the Population
To tackle the binary data, we take a statistical perspective in this paper. Several studies show many prevalent diseases and common medical conditions are often associated with commonalities among populations, such as shared living environments or behavioral habits [1]. In terms of epidemiology, temporal specificity and spatial specificity inevitably lead to a closer medical-relatedness among patients within the same sub-cohort compared to patients belonging to other sub-cohorts [26, 1, 19]. Inspired by this observation, we make two assumptions:
Assumption 1
the outcome of a medical event $E_{i}$ of the entire population is subject to a Bernoulli distribution.
Assumption 2
an EHR dataset is an i.i.d. sample from the population.
As such, an intuitive solution to the signal-stopping 0 entries is to represent them with the Bernoulli mean of the corresponding medical event $1-P(E_{i}=1)$. This Bernoulli mean can be empirically estimated from the EHR dataset which is assumed to be subject to the identical distribution as the population.
By mapping the 0-1 events to Bernoulli statistics, the equally important Yes/No outcomes have continuous representations. This can be important since the problem of data sparsity has been one of the most prominent obstacles to effectively mining the EHR data, as zero entries in practice do not propagate any learning signal. For instance, in deep learning models, a value of 0 cannot effectively activate the connection weights within a projection layer. Similarly, in graph-based models, zero values have no contribution to node attributes and message passing between nodes. As Bernoulli’s mean can take on any value between 0 and 1, this transformation also allows for more effective representations of the patient-event/event-event relationship. Quantifying such a relationship helps reflect the intrinsic connection between medical events, and contributes more to accurate medication recommendation than binary signals [16].
Figure 2: (Left) We transform the $\{0,1\}$ entries into Bernoulli means empirically estimated from the EHR samples, with the assumptions detailed in Section 4.1. (Middle) Graph learning directly on $\{0,1\}$ data can stall learning since zero entries prevent learning signals from propagating. (Right) Our proposed method learns from the Bernoulli transformed entries. Here, edges $e_{ij}$ are constructed from Bernoulli means, and nodes $x_{i}$ are constructed from conditional Bernoulli probabilities, see Algorithm 1.
@@ 4.1 Patient-Event Graph Construction
We propose to learn by graph neural networks (GNNs) the inherent relationship between medical events within each patient sample as well as between an event and the patient population. A system overview has been shown in Fig. 2. Specifically, for each patient a graph $\mathcal{G}_{n}=(\mathcal{V}_{n},\mathcal{E}_{n})$ is constructed, where $\mathcal{V}_{n}$ denote the set of vertices, corresponding to the patient-event relationship; $\mathcal{E}_{n}$ the set of edges representing the event-event relationship. If we naïvely initialized the $\mathcal{V}_{n},\mathcal{E}_{n}$ by the elements of $\mathbf{X}$, the zero entries may cause significant loss of information of the non-occurrence of events since zero values typically stall the propagation of learning signals. Therefore, we propose to initialize them in the following approach.
@@
Patient-to-patient correlation embedding. By assumption (i) in Section 3.2, the occurrence of an event is subject to a Bernoulli distribution $P(E_{j}=k)=\rho^{k}(1-\rho)^{k}$, where $k\in\{0,1\}$, and $\rho$ denotes the mean of the distribution. By assumption (ii), the event in the EHR data sample is subject to the same distribution $P(\mathbf{X}_{\cdot j}=k)=\rho^{k}(1-\rho)^{k}$, where we denote the $\mathbf{X}_{\cdot j}$ denote the event $E_{j}$ for all patient samples. Therefore, we propose to initialize the vertices by the Bernoulli mean $\rho$. While this quantity is not known, we can replace it with a sample estimate:
$\tilde{\rho}:=\tilde{P}(X_{\cdot j}=1)=\frac{\sum_{i=1}^{N}\mathbf{X}_{ij}}{N},\,\,\tilde{P}(X_{\cdot j}=0)=1-\tilde{\rho}.$
@@
Event-to-event correlation embedding. Following this idea, we propose to initialize the edges representing the event-to-event relationship by conditional Bernoulli probability. Intuitively, an edge from event $j$ to event $i$ encodes some information conditional on the observation of $i$ [4, 17]. From this consideration, we initialize the edge $e_{ij}$ from event $E_{j}$ to event $E_{i}$ with:
$e_{ij}:=P(E_{i}=1|E_{j}=1)=\frac{{P}(E_{i}=1\text{ and }E_{j}=1)}{{P}(E_{j}=1)}.$
Again, while the exact distribution is not known, we can replace it with sample estimate: the conditional probability is instead estimated by
$\frac{P(\mathbf{X}_{\cdot i}=1\text{ and }\mathbf{X}_{\cdot j}=1)}{P(\mathbf{X}_{\cdot j}=1)}=\frac{\#\{\text{patients with positive event }i\text{ and }j\}}{\#\{\text{patients with positive event }j\}}.$
It is worth noting that the proposed initialization can be seen as a transformation from binary data to continuous values. In the experiments, we shall also compare against other transformation methods.
Algorithm 1 Pseudocode of patient-event graph construction.

### v3

@@ Abstract
Electronic Health Records (EHRs) frequently contain binary event sequences, such as diagnosis, procedure, and medication histories. These sequences, however, are characterized by a large percentage of zero values, presenting significant challenges for model learning due to the scarcity of gradient updates during processing. We propose a novel approach that reframes EHR data from a statistical perspective, treating it as a sample drawn from patient cohorts and transforming binary events into continuous Bernoulli probabilities. This transformation not only models deterministic binary events as distributions but also captures event-event relationships through conditional probabilities. Building upon this probabilistic foundation, we implement a graph neural network that learns to capture complex event-event correlations while emphasizing patient-specific features. Our framework is evaluated on the task of medication recommendation, demonstrating state-of-the-art performance on large-scale databases. Extensive experimental results show that our proposed BernGraph  [^1] The source code is available at https://anonymous.4open.science/r/BEHRMecom-031C achieves 0.589, 0.846, and 0.844 of Jaccard, F1 score, and PRAUC, respectively, on the MIMIC-III dataset, with improvements of 4.6%, 6.7%, and 6.3% compared to baselines, including those using secondary information. This work presents a promising direction for handling sparse binary sequences in EHRs, potentially improving various healthcare prediction tasks.
@@ 1 Introduction
Medication errors are one of the most serious medical errors that could threaten patients’ lives. Studies indicate that over 42% of these errors stem from physicians or doctors with insufficient experience or knowledge concerning specific drugs and diseases (Tran et al., 2021). This challenge becomes even more pronounced with more new drugs and treatment guidelines. On the other hand, many patients are concurrently diagnosed with multiple diseases, complicating the decision-making process for doctors when determining more appropriate remedies for patients (Wu et al., 2022). To this end, the medical community will benefit from an automated medication recommendation system that can assist doctors in decision-making. In the past decades, a vast amount of clinical data representing patient health status (e.g., medical reports, radiology images, and allergies) has been collected. This has remarkably increased digital information, known as electronic health records (EHR), available for patient-oriented decision-making. In this context, mining EHR data for medication recommendation has attracted growing research interests (Wu et al., 2022), as illustrated in Fig. 1(a).
@@
Successful methods typically rely on modeling outcomes of various medical events and diagnoses by observing longitudinal patient histories, i.e., tracking state variations across hospital visits (Choi et al., 2016a; Zhang et al., 2017; Yang et al., 2021b; Wu et al., 2022). They often employ sequential neural networks, such as RNNs or Transformers. They construct a feature space centered around visits, each characterized by the current state of various medical events. Their objective is to discern long-range temporal dependencies that reflect the historical progression of diseases. These methods, however, suffer from a number of drawbacks. From a data viewpoint, different EHR databases adopt distinct visit recording methodologies (Singh et al., 2015). For example, the AMR-URI database (Michael et al., 2020) employs a regular visit record, while visits in the MIMIC database (Johnson et al., 2016) are patient-specific. Even within the same database, the time intervals between visits can differ greatly between patients. From a model perspective, a common obstacle is the "cold-start" issue, where medication recommendation systems struggle due to insufficient patient-specific historical data (Ye et al., 2021). In practice, this issue becomes particularly pronounced when the system encounters newly onboarded patients with scarce or even absent historical data (i.e., visits) about their medical backgrounds. While many researchers treat irregular time intervals as input variables (Nguyen et al., 2016; Miotto et al., 2016; Xiang et al., 2019) or apply the carry-forward imputation method (Luo, 2022), these approaches may introduce biases to temporal dependencies.
Figure 1: (a) Medication recommendations based on historical medical events. (b) The vast majority of event data are zero values. For example, the proportion of non-zero values exceeding 10% is lower than 0.9% in MIMIC-III. (c) Those zero values do not propagate learning signals.
@@
Some studies focus solely on current diagnoses of patients and medical event records, modeling event-event correlations for recommendation purposes without considering historical visits (Zhang et al., 2017; Su et al., 2020). However, a performance gap remains when compared to visit-based methods. A typical medical event has a binary value indicating its outcome: 1 for occurrence and 0 for non-occurrence, which leads to a sparsity issue in EHR data. The sparsity, characterized by a high proportion of zero entries in event sequences, can significantly impact the performance of recommendations (seen in Fig. 1(b)). Zero values provide no tangible information for deep learning models, preventing the activation of weight propagation. This often leads to a vanishing gradient problem and results in collapsed representations for medication recommendations (as shown in Fig. 1(c)). Previous studies employ secondary information, e.g., Drug-Drug Interaction Knowledge Graph (DDI-KG), to complement information to medical events (Wu et al., 2022). Yet, sparsity remains a persistent issue.
@@
Therefore, this paper presents BernGraph, a binary EHR data-oriented drug recommendation system. Our method is built upon a graph neural network (GNN) and solely utilizes the EHR binary medical events for recommendation. To tackle the sparsity in binary event sequences, we take a statistical perspective that allows us to transform the $(0,1)$ values into continuous probabilistic representations and estimate them from patient cohorts. This strategy has demonstrated substantial efficacy in modeling graphical structures and correlations within binary EHR data. Architecturally, diverse events are incorporated as node attributes within the GNN. When two events co-occur, we model their relationship as conditional Bernoulli probabilities to initialize the edge between them. This allows GNN to learn the correlations between events across different patients for medication recommendations. Our results demonstrate that our method outperforms several benchmarks, including the latest state-of-the-art (SOTA) visit-based methods. Specifically, our method shows an improvement on the MIMIC III dataset of approximately 4.8%, 6.7%, 6.3%, and 4% in terms of Jaccard, F1-score, PRAUC, and AUROC, respectively. Moreover, BernGraph is a simple and effective model that complements existing research in which the EHR data is often exploited in conjunction with secondary data but without transformations.
@@ 2 Related Works
- EHR-based Medication Recommendation. Historical Visit-based Methods: Many studies leverage deep learning to extract features from EHR visit variations (Choi et al., 2016b; Wang et al., 2018; W et al., 2019; Bhave and Perotte, 2021; Su et al., 2020). For example, DoctorAI (Choi et al., 2016b) employs RNNs to capture dependencies between patient histories and doctor recommendations. MICRON (Yang et al., 2021a) focuses on medication changes and combinations from the last visit. Some works (Yang et al., 2021b; Wu et al., 2022) further incorporate additional information like Drug-Drug Interactions (DDI) to aid the model predictions.
Current Medical Event-based Methods: On the other hand, some works aim to analyze the patient’s current health status: recommendations are made based on the EHR records from each patient visit (Zhang et al., 2017; Gong et al., 2021; Shu et al., 2023). LEAP (Zhang et al., 2017) encodes patient information in a one-hot encoding and then learns the relationship between patient features and drugs. SMR (Gong et al., 2021) combines EHR data with an external medical knowledge graph to leverage external medical expertise in decision-making. MT-GIN (Shu et al., 2023) models event-event correlations in EHR data using a graph-based approach, with a focus on events that co-occur. This approach overlooks the information that 0 values in EHR binary data can potentially provide.
- Graph Modeling for Medication Recommendation. Recent works leverage graph structures to model event correlations (Choi et al., 2016b, 2020; Su et al., 2020; Wharrie et al., 2023; Shang et al., 2019; Gong et al., 2021; Tan et al., 2022; Wu et al., 2022; Shu et al., 2023). Gamenet (Shang et al., 2019) and Cognet (Wu et al., 2022) use GNNs to capture correlations, integrating DDI data for representation learning. GATE (Su et al., 2020) proposes a GNN-based approach using events from a single visit. COGNet (Wu et al., 2022) further integrates both EHR and DDI graphs into a generative model. However, these approaches are designed based on binary medical event data.
- Binary Data Representation. Mining binary data is challenging due to the varied interpretations of 0s and 1s across contexts (D and Stepniewska, 1999; Verstrepen et al., 2017). Additionally, 0s often contribute little to deep learning. In the EHR context, medGAN and POPCORN transform binary events into continuous probabilities (Choi et al., 2017; Bhave and Perotte, 2021), but these focus on historical data or overall distributions. Developing an effective method to represent binary events directly in EHR data remains challenging.
Figure 2: (a) We transform $\{0,1\}$ EHR entries into Bernoulli means. (b) Learning directly from $\{0,1\}$ data hinders signal propagation. (c) Our method learns from continuous entries, with edges constructed from Bernoulli means and nodes from conditional Bernoulli probabilities (see Algorithm 1).
@@ 3.1 Recommendation System Based on EHR
EHR data consists of various medical events, such as multiple symptoms, recorded chronically from patients. Each data sample contains binary outcomes of these events. Let $\boldsymbol{x}\in\{0,1\}^{M}$ denote a patient EHR sample, where $M$ represents the total number of medical events, $1$ stands for the occurrence of the corresponding event, and 00 for the absence. The collection of EHRs is a matrix $\mathbf{X}\in\{0,1\}^{N\times M}$ composed by stacking $N$ patient vectors $\{\boldsymbol{x}\}_{1:N}$. It is challenging to build effective recommendation systems using solely the EHR data, mainly due to the following reasons: (1) binary data poses a challenge to effective machine learning: while both 0 and 1 play an equally important role in identifying a proper recommendation, in practice, 0 entries generally do not propagate learning signals; (2) medical events differ across platforms, and as a result, two binary matrices may share the same shape but have entirely different underlying meanings. Our goal is to mine the complete information hidden in binary EHR data effectively, based on which an effective, general-purpose recommendation system can be built. Formally, BernGraph can be described by $F(\cdot):\{0,1\}^{M}\rightarrow\{0,1\}^{C}$, which accepts an EHR sample as input, and outputs $C$ binary indicators indicating whether or not recommend drug.
@@ 3.2 EHR as a Sample from the Population
This paper introduces a statistical method to address binary data. Several studies suggest prevalent diseases and medical conditions often share commonalities across populations due to factors like environment and behavior (Bache et al., 2013). In terms of epidemiology, temporal and spatial specificity lead to closer medical relatedness within sub-cohorts (Sun et al., 2012; Bache et al., 2013; Poulakis et al., 2022). Based on this, we assume that (i)the outcome of a medical event $E_{i}$ follows a Bernoulli distribution for the population, and (ii)the EHR dataset is an i.i.d. sample from this population.
To address the issue of signal-stopping 0 entries, we represent them using the Bernoulli mean $1-P(E_{i}=1)$, empirically estimated from the EHR dataset, which assumes the same distribution as the population. This transforms binary 0-1 events into continuous representations, helping mitigate the problem of data sparsity. In deep learning models, a value of 0 prevents activation of connection weights, and in graph-based models, zeros do not contribute to node attributes or message passing. The Bernoulli mean, ranging from 0 to 1, provides more effective representations of patient-event and event-event relationships, leading to more accurate medication recommendations than binary signals (M.M. et al., 2015).
@@ 4.1 Patient-Event Graph Construction
We propose using GNNs to learn the relationships between medical events within each patient and across the patient population. Fig. 2 provides a system overview. For each patient, we construct a graph $\mathcal{G}_{n}=(\mathcal{V}_{n},\mathcal{E}_{n})$, where $\mathcal{V}_{n}$ represents patient-event relationships and $\mathcal{E}_{n}$ captures event-event connections. Naïvely initializing $\mathcal{V}_{n}$ and $\mathcal{E}_{n}$ from $\mathbf{X}$ could lead to information loss from zero entries, as they hinder signal propagation. To this end, we propose an alternative initialization method for graph construction.
@@
Patient-to-patient correlation embedding. By assumption(i) in Section 3.2, the occurrence of an event is subject to a Bernoulli distribution $P(E_{j}=k)=\rho^{k}(1-\rho)^{k}$, where $k\in\{0,1\}$, and $\rho$ denotes the mean of the distribution. By assumption(ii), the event in the EHR data sample is subject to the same distribution $P(\mathbf{X}_{\cdot j}=k)=\rho^{k}(1-\rho)^{k}$, where we denote the $\mathbf{X}_{\cdot j}$ denote the event $E_{j}$ for all patient samples. Therefore, we propose initializing the vertices by the Bernoulli mean $\rho$. While this quantity is not known, we can replace it with a sample estimate:
$\tilde{\rho}:=\tilde{P}(X_{\cdot j}=1)=\frac{\sum_{i=1}^{N}\mathbf{X}_{ij}}{N},\,\,\tilde{P}(X_{\cdot j}=0)=1-\tilde{\rho}$ (1)
@@
Event-to-event correlation embedding. Then, we propose to initialize the edges representing the event-to-event relationship using conditional Bernoulli probability. Intuitively, an edge from event $j$ to event $i$ encodes some information conditional on the observation of $i$ (Choi et al., 2020; Murphy, 2023). From this consideration, we initialize the edge $e_{ij}$ from event $E_{j}$ to event $E_{i}$ with:
$e_{ij}:=P(E_{i}=1|E_{j}=1)=\frac{{P}(E_{i}=1\text{ and }E_{j}=1)}{{P}(E_{j}=1)}$ (2)
Since the exact distribution is not known, we hence replace it with a sample estimate. That is, the conditional probability is instead estimated by $\frac{P(\mathbf{X}_{\cdot i}=1\text{ and }\mathbf{X}_{\cdot j}=1)}{P(\mathbf{X}_{\cdot j}=1)}$ where $P(\mathbf{X}_{\cdot i}=1\text{ and }\mathbf{X}_{\cdot j}=1)$ is the ratio of patients with both events $i$ and $j$ being positive, and $P(\mathbf{X}_{\cdot j}=1)$ is the ratio of patients with event $j$ being positive. It is worth noting that the proposed initialization can be seen as a transformation from binary data to continuous values.
4.2 Patient-to-Event Graph Learning
In Section 4.1, we mapped each patient sample $\mathbf{X}_{i\cdot}$ to a graph $\mathcal{G}_{i}=(\mathcal{V}_{i},\mathcal{E}_{i})$, with empirical Bernoulli mean initialization. This transformation converts the original drug recommendation binary classification problem into a graph classification problem. Specifically, we first map the EHR data into a set of vertices and edges $\{0,1\}^{M}\stackrel{{\scriptstyle\text{B.I.}}}{{\rightarrow}}\mathcal{V}_{i},\mathcal{E}_{i}\in[0,1]^{M}$, where B.I. denote Bernoulli Initialization. A graph learning function is then applied to gain drug recommendation outputs $F(\cdot):\mathcal{V}_{i},\mathcal{E}_{i}\rightarrow\{0,1\}^{C}$. Notably, we employ EGraphSage (Lo et al., 2022) as our $F$, where the GNN learns the node and edge embeddings $h_{\nu},h_{e}$ by an MLP. The first layer has an interpretation of nonlinearly transforming the Bernoulli statistics residing on the graph we initialized. Graph learning aims to discern correlations between events through message aggregation and learn the event (node) attributes. The message $m_{j}$ on each edge $e_{ij}$ from node $i$ to node $j$ is:
$m_{j}=\text{relu}\left(\mathbf{W}_{m}\times\left(h_{\nu_{j}}\oplus h_{e_{ij}}\right)\right)$ (3)
where $\oplus$ represents concatenation. $\mathbf{W}_{m}$ is a learnable weight matrix for message passing, and relu is a non-linear activation function.
Message aggregation: Afterward, we aggregate messages $m_{j}$ across edges for each node:
$\text{AGGREGATE}_{i}=\frac{1}{|\mathcal{N}(i)|}\sum_{j\in\mathcal{N}(i)}m_{j}$ (4)
where $\mathcal{N}(i)$ represents the set of neighbors.
$h_{\nu_{i}}=\text{activation}\left(\mathbf{W}_{u}\times\left(h_{\nu_{i}}\oplus\text{AGGREGATE}_{i}\right)\right)$ (5)
where $h_{\nu_{i}}^{\prime}=\frac{h_{\nu_{i}}^{\prime}}{\|h_{\nu_{i}}^{\prime}\|_{2}}$ is the updated feature of node $i$, and $\mathbf{W}_{u}$ is learnable weight matrix for updating, .$\|.\|_{2}$ denotes the L2 norm.
Algorithm 1 Pseudocode of patient-event graph